OpenStudy (anonymous):

Find sin (a/2) if cos(a)= 12/13 for 0 degrees≤ a≤ 90 degrees

OpenStudy (jdoe0001):

hmmm if the angle is between \(\bf 0^o < a < 90^o\ that'd be 1st Quadrant what signs are sine and cosine? is sine + or -? is cosine + or - anyway?

OpenStudy (jdoe0001):

hmmm well.... just notice.. you just need the half-angle... hmmm ok

OpenStudy (anonymous):

It does't say so i assume they are both positive

OpenStudy (jdoe0001):

well just bear in mind that -> \(\bf sin\left(\cfrac{{\color{red}{ \theta}}}{2}\right)=\pm \sqrt{\cfrac{1-cos({\color{red}{ \theta}})}{2}}\) and you're given the cosine already keep in mind that you're in the 1st Quadrant, so sine has a sign there, and that's the root you'd keep

OpenStudy (anonymous):

If it's in the first quadrant, that means its positive, right?

OpenStudy (jdoe0001):

yeap, which means \(\bf sin\left(\cfrac{{\color{red}{ \theta}}}{2}\right)=+ \sqrt{\cfrac{1-cos({\color{red}{ \theta}})}{2}}\)

OpenStudy (anonymous):

So how do i find theta?

OpenStudy (jdoe0001):

well.. you're asked to find sin(a/2) you're already given the cosine of angle "a" anyhow and that's all you'd need to get sin(a/2)

OpenStudy (anonymous):

So all i have to do is input 12/13 for theta into the equation and solve?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

yeap pretty much, yes

OpenStudy (perl):

do you need help

OpenStudy (anonymous):

I got -6, but thats not an option for me

OpenStudy (jdoe0001):

\(\bf sin\left(\cfrac{{\color{red}{ \theta}}}{2}\right)=+ \sqrt{\cfrac{1-\left({\color{brown}{ \frac{12}{13}}}\right)}{2}}\)

OpenStudy (jdoe0001):

-6? from a square root...... ?

OpenStudy (jdoe0001):

keep in mind that 1 = 13/13 thus one could say \(\large { sin\left(\cfrac{{\color{red}{ a}}}{2}\right)=+ \sqrt{\cfrac{\frac{13}{13}-\left({\color{brown}{ \frac{12}{13}}}\right)}{2}} }\)

OpenStudy (anonymous):

On the top where I have 12/13, wouldn't i have to get a common denominator?

OpenStudy (jdoe0001):

yues

OpenStudy (jdoe0001):

anyhow.. yes.. but doesn't end up with -6 though

OpenStudy (anonymous):

then I'm left with 1/13

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

and then divide that by 2 and find the square root?

OpenStudy (anonymous):

I got it! thank you so much(:

OpenStudy (jdoe0001):

yeap

OpenStudy (jdoe0001):

yw